OpenStudy (anonymous):

8y - 3x = 6 solve for y

OpenStudy (anonymous):

add -3x on both sides and get 8y = 6+3x divide 8 on both sides y = (6+3x)/8 simplify y = 3x/8 + 3/4

OpenStudy (anonymous):

Thank you